# STG-Mamba: Spatial-Temporal Graph Learning via Selective State Space Model

### v2

@@ 3.3 GS3B: Graph Selective State Space Blocks
Figure 2: The selection mechanism of STG-Mamba that facilitates input-dependent feature selection and focusing.
Inspired from Mamba, the GS3B is the major component of STG-Mamba, as depicted in Figure 1. After being processed by Layer Normalization, the input is divided into two sub-branches. In the first branch, the input STG embedding passed through a Linear Layer, followed by the proposed KFGN network and an activation function, before entering the GSS-Layer for selective feature learning. After that, the learned features are normalized by the Layer Norm. For the second branch, the input STG embedding undergoes a Linear Layer and an activation function. Subsequently, the two branches are fused together by an element-wise dot product operation. Finally, the merged embedding passes through a linear layer, and the outcome is combined with a residual connection to form the GS3B block’s output. In this research, Sigmoid Linear Unit (SiLU) is employed as the default activation function. We further comprehend the detailed process of the STG-Mamba’s Selection Mechanism in Figure 2.
The GSS-Layer is a graph selective state space module specifically designed for STG data. GSS-Layer incorporates a selective mechanism which is capable of adjusting the SSM model’s parameters according to the input features. The characteristics enable the model to distinguish and retain key information which contributes significantly in STG learning tasks, while filtering out the irrelevant information. The pseudo-code of GSS-Layer (graph selective state space algorithm) is elaborated in Algorithm 1.
Algorithm 1 Pseudo Code of the Proposed Graph Selective State Space Algorithm
Input: $X\in\mathbb{R}^{batch\times L\times dim}$, the historical ground-truth STG data, where $batch$ is batch size of model training, $L$ is the historical time window length, $dim$ is the feature dimension of STG data;
$\mathbb{\alpha_{t}}$, the graph adjacency matrix of the current time step calculated by Deep Kalman Filter Network.
Params: $\mathbb{A}$ is the network param: nn.Parameter; $\mathbb{D}$ is the network param: nn.Parameter;
Operator: $Linear(\cdot)$, the linear projection layer;
Output:$y$, the processed feature with shape $\mathbb{R}^{batch\times L\times dim}$;
1:  $\Delta,\mathbb{B},\mathbb{C}=Linear(X),Linear(X),Linear(X)$
2:  $\overline{\mathbb{A}}=\exp(\Delta\mathbb{A})$
3:  $\overline{\mathbb{B}}=(\Delta\mathbb{A})^{-1}(\exp(\Delta\mathbb{A})-I)\cdot\Delta B$
4:  $h_{t}=(\overline{\mathbb{A}}\odot\mathbb{\alpha_{t}})h_{t-1}+\overline{\mathbb{B}}x_{t}$
5:  $y_{t}=\mathbb{C}h_{t}+\mathbb{D}x_{t}$
6:  $y=[y_{1},y_{2},...,y_{t},...,y_{L}]$.
7:  return $y$
@@ 3.4 Computational Efficiency of STG-Mamba
GS3B’s selection mechanism is introduced in Algorithm 1, which outlines the graph selective state space algorithm for efficiently processing STG data. It starts with initializing linear projections of the input STG data to obtain parameters $\Delta,\mathbb{B},\mathbb{C}$, which are used to compute the modified state transition matrix $\overline{\mathbb{A}}$ and input matrix $\overline{\mathbb{B}}$. These matrices incorporate the graph adjacency matrix $\mathbb{\alpha_{t}}$, updated by a Deep Kalman Filter Network, to calculate the hidden state $h_{t}$ for each time step. Finally, the output $y_{t}$ is generated by applying linear transformations to the hidden state and the input data, aggregating the results over the historical time window to form the output feature $y$. This algorithm effectively integrates dynamic graph structure information into SSSM, enhancing its capability to model sophisticated STG dynamics.
@@ 4.1 Dataset Statistics and Baseline Methods
Datasets. For experimental evaluation, we meticulously select three real-world STG datasets from California road network speed records (PeMS04) Song et al. (2020), Hangzhou urban metro system entry/exit records (HZMetro) Liu et al. (2022), and weather station records across mainland China (KnowAir) Wang et al. (2020b). Acquired from the Caltrans Performance Measurement System (PeMS), PeMS04 records the average traffic speed from a total number of 307 sensors distributed on California Highway system. The time range of PeMS04 is 1st/Jan/2018 - 28th/Feb/2018, with 5-minute time interval. HZMetro was created with the In/Out transaction records of 80 stations of Hangzhou Metro system from 1st Jan 2019 to 25th Jan 2019, which takes 15-minute as data collection time interval. KnowAir documents weather observations in every 3-hour, encompassing a dataset that spans from 2015 to 2018. Within the KnowAir dataset, there are three subsets. In this study, we exclusively use subset No. 3, covering the period from 1/Sep/2016 to 31/Jan/2017. The detailed dataset statistics and descriptions are summarized in Table 1.
@@
STGNN-based Methods:
(I) STGCN Yu et al. (2018): STGCN propose to capture the complex spatial-temporal dependencies within traffic network by graph neural networks, and design a fully convolutional approach for fast training.
(II) STSGCN Song et al. (2020): STSGCN presents a novel model which efficiently handling complex localized spatial-temporal correlations and data heterogeneity through a synchronous mechanism.
(III) STG-NCDE Choi et al. (2022): STG-NCDE leverages the concept of Neural Controlled Differential Equations (NCDEs) for spatial-temporal processing within a unified framework, achieving superior accuracy across six benchmark datasets.
(IV) DDGCRN Weng et al. (2023): DDGCRN dynamically captures the spatial-temporal features by separating the normal and abnormal signals and employing a novel graph convolutional recurrent network approach.
Attention (Transformer)-based Methods:
(V) ASTGCN Guo et al. (2019): ASTGCN leverages an attention-based spatial-temporal graph convolutional network to accurately predict STG data by dynamically capturing complex spatiotemporal correlations and patterns.
(VI) ASTGNN Guo et al. (2021): Attention based Spatial-Temporal Graph Neural Network incorporates a unique trend-aware self-attention and dynamic graph convolution to address temporal dynamics, spatial correlations, and the inherent periodicity and heterogeneity of STG data.
(VII) PDFormer Jiang et al. (2023): PDFormer introduces a tailored Transformer model, adept at capturing dynamic spatial dependencies and addressing the time delay in traffic condition propagation with its unique spatial self-attention and traffic delay-aware feature transformation.
(VIII) STAEformer Liu et al. (2023): STAEformer introduces a novel approach that enhances vanilla Transformer with a spatiotemporal adaptive embedding component, enabling it to capture complex STG patterns.
(IX) MultiSPANS Zou et al. (2024): MultiSPANS leverages multi-filter convolution modules, Transformers, and structural entropy optimization to effectively model complex multi-range dependencies in STG forecasting.
@@ 4.2 Implementation Settings
We split all the datasets with a ratio of 6:2:2 for Training/Validation/Testing sets along the time axis, respectively. Before starting model training session, all the data samples are normalized into range $[0,1]$ with MinMax normalization method. We set the batch size as 32 and employ AdamW as the Optimizer. The learning rate is initialized as $1e^{-3}$, with a weight decay of $5e^{-2}$. CosineAnnealingLR Loshchilov and Hutter (2016) is adopted as the learning rate scheduler, with a maximum of 50 iterations, and a minimum learning rate $1e^{-5}$. Model training epochs is set as 300. The number of Encoder/Decoder Layer is set as $N=4$ and $N^{\prime}=4$, respectively. We use the historical ground-truth data from the past 12 steps to forecast one future time step. Mean Squared Error (MSE) is employed as the Loss Function for optimization. And finally, we employ RMSE, MAE, MAPE as the evaluation metrics. All experiments are conducted under one NVIDIA A100 80GB GPU for neural computing acceleration.
@@ 4.3 Results Evaluation and Comparison
4.3.1 Experimental Results Comparison on Three STG Forecasting Tasks
@@
4.3.2 Computational Efficiency Evaluation
STG-Mamba offers substantial improvement in computational efficiency in addition to performance gain. In order to quantatively evaluate computational efficiency, we select the widely recognized Floating Point Operations (FLOPs) and Inference Time as criteria. For both of the two criteria, a smaller value indicates better model performance. STG-Mamba is compare with other Transformer-based benchmark methods: ASTGNN, PDFormer, and STAEformer, on the same testing set.
Table 3 showcases the inference time comparison on PeMS04 and KnowAir testing set. It can be seen that ASTGNN is relatively redundancy, with the lowest inference time among the compared Transformer-based methods. STG-Mamba gains 18.09% and 20.54% performance increase in inference time than the 2nd best performance model (STAEformer).
We further compare the computational FLOPs increment with accordance to the scale complexity of STG network. Specifically, we use the number of node individuals to represent the scale of STG network. In the experiments, we select different number of nodes from PeMS04 and KnowAir. We select $[50,100,150,200,250,300]$ and $[30,60,90,120,150,180]$ node individuals for PeMS04 and KnowAir dataset, respectively. STG-Mamba is compared with the benchmark STAEformer in every settings. Figure 4 visualizes the FLOPs evaluation under different STG scale complexity. We can easily find that STG-Mamba enjoys a linear increment in FLOPs, while STAEformer shows quadratic increment in FLOPs. At the initial small-scale setting, the FLOPs for STG-Mamba/STAEformer are quite similar. However, with the increasing of network complexity and scale, the difference becomes more significant. At the 300 node setting of PeMS04, the FLOPs for STG-Mamba and STAEformer is 36.85G and 92.49G (2.50 times more), respectively. At the 180 node setting of KnowAir, the FLOPs for STG-Mamba and STAEformer is 57.08G and 118.13G (2.07 times more), respectively.
Benefiting from the linear computational increment brought by modern selective state space models, STG-Mamba and alike models demonstrate stronger capability and computational efficiency for modeling large-scale STG data, which holds promise for addressing the computational overhead in future large-scale STG learning.
Table 3: Test inference time comparison with benchmark Transformer-based approaches.
Model Inference Time on PeMS04 Inference Time on KnowAir
ASTGNN 4.85 s 5.31s
PDFormer 3.93 s 4.28s
STAEformer 3.76 s 4.09s
STG-Mamba 3.08 s 3.25s
performance $\uparrow$ +18.08% +20.54%
Figure 4: (Computational Efficiency Eval) FLOPs comparison between STG-Mamba and STAEformer on PeMS04/KnowAir dataset under different STG node number settings.
@@ 4.5 Ablation Study
    STG-Mamba w/o KFGN: We replace the proposed Kalman Filtering Graph Neural Networks (KFGN) within STG-Mamba with the standard Graph Convolution Network (GCN).
@@
    STG-Mamba w/o GSS-Layer: The GSS-Layer with proposed Graph Selective State Space Algorithm is replaced by the standard selective state space module as introduced in Gu and Dao (2023).
@@
    Transformer w KFGN: The GS3B block is reformulated as Transformer block (similar to ASTGNN Guo et al. (2021)), which consists of attention mechanism and the proposed KFGN.
As illustrated in Figure 5, STG-Mamba w/o KFGN shows slight worse performance than the full model, demonstrating the necessity and suitability of employing Kalman filtering-based dynamic GNN for SSSM-based methods. Different from the mainstream methods which employ various attention mechanisms to upgrade STG network’s adjacency matrix, our KFGN is more consistent with the SSSM context, and at the same time keep the linear algorithm complexity. STG-Mamba w/o GSS-Layer performs even worse than w/o KFGN. This is because in our design, GSS-Layer will take the generated STG adjacency matrix $\alpha_{t}$ at this timestamp as input for selective state space algorithm’s better upgradation. Replacing GSS-Layer with the standard selective state space layer leads to a less suitable version for STG learning. Finally, when we change from Mamba-based architecture to Transformer-based architecture, there is a substantial performance drop among all the three tasks. This discovery proves the effectiveness of modern SSSM over attention-based designs.
@@ 5 Conclusions
In this work, we first propose a novel spatial-temporal graph (STG) learning model named STG-Mamba that leverages modern selective state space models (SSSM) for accurate and efficient STG forecasting tasks. STG-Mamba is formulated as Encoder-Decoder structure, with the Graph Selective State Space Block (GS3B) as basic module. GS3B’s selection method facilitates input-dependent edge construction and node feature selection, successfully integrates STG network with SSSMs. Through the Kalman Filtering Graph Neural Networks (KFGN), the STG network achieves a smooth and dynamic evolution process of spatiotemporal graph adjacency matrix upgrading, which align with the whole graph selective state space modeling process. Compared with Attention-based methods, our STG-Mamba is able to achieve linear-time complexity and substantial reduction in FLOPs and test inference time. We conduct extensive empirical studies on three open-sourced STG forecasting datasets from urban traffic, metro In/Out flow, and weather records. Experimental results demonstrate the consistent superiority of STG-Mamba over other benchmark methods in both forecasting performance and computational efficiency. We believe that STG-Mamba presents a brand new promising approach to general STG learning fields, offering competitive model performance and input-dependent context-aware feature selection under an affordable computational cost.

### v3

@@ 3.3 ST-S3M: Spatial-Temporal Selective State Space Module
Figure 3: Illustration of the spatial-temporal state space selection mechanism that facilitates input-dependent feature selection and focusing.
ST-S3M plays the significant role of adaptive STG feature selection, serving as the counterpart of Attention mechanism. The architecture of ST-S3M is shown in Figure 1. We also give an vivid illustration of the STG feature selection procedure in Figure 4.
After being processed by KFGN module for dynamic spatiotemporal dependency modeling and statistical-based integration & upgrading. The generated embedding $\hat{h}_{fuse}$ is sent to ST-S3M for input-specific dynamic feature selection. In ST-S3M, the input STG embedding passes through a Linear Layer, followed by a splitting operation. Let $b$ denotes batch size, $l$ denotes sequence length, $d_{model}$ be the feature dimension, $d_{inner}$ be the model inner feature dimension, we have:
$h_{\text{main-res}}$ $=W_{in}\hat{h}_{fuse}+b_{in}$
$(h_{\text{main}},res)$ $=\text{split}(h_{\text{main-res}})$ (14)
where $\hat{h}_{fuse}\in\mathbb{R}^{b\times l\times d_{model}}$, $W_{in}\in\mathbb{R}^{2d_{inner}\times d_{model}}$, $b_{in}\in\mathbb{R}^{2d_{inner}}$, $h_{\text{main-res}}\in\mathbb{R}^{b\times l\times 2d_{inner}}$, $h_{\text{main}},res\in\mathbb{R}^{b\times l\times d_{inner}}$
Then, $h_{\text{main}}$ flows into a 1D convolution layer, followed by SiLU activation function:
$h_{\text{main}}^{{}^{\prime}}=\text{Conv1D}(h_{\text{main}})$
$h_{\text{main}}^{{}^{\prime\prime}}=\text{SiLU}(h_{\text{main}}^{{}^{\prime}})$ (15)
where $h_{\text{main}}^{{}^{\prime\prime}}\in\mathbb{R}^{b\times l\times d_{inner}}$. The output of SiLU is sent to state space selection mechanism for processing:
$h_{\text{sssm}}=\text{SSSM}(h_{main}^{{}^{\prime\prime}})$ (16)
where $h_{\text{sssm}}\in\mathbb{R}^{b\times l\times d_{inner}}$. On the same time, the residual part $res$ is also processed by SiLU activation, and finally we employ element-wise dot product to fuse the main STG embedding and its residual: $h_{\text{sssm}}\odot\text{SiLU}(res)$. The fused result is transformed through a Linear projection:
$h_{out}=W_{out}(h_{\text{sssm}}\odot\text{SiLU}(res))+b_{out}$ (17)
where $h_{out}\in\mathbb{R}^{b\times l\times d_{model}}$ is the final output of ST-S3M. $W_{out}\in\mathbb{R}^{d_{model}\times d_{inner}}$, $b_{out}\in\mathbb{R}^{d_{model}}$.
The State Space Selection Mechanism in ST-S3M plays a pivotal role in adaptive spatiotemporal feature selection. We detail the parameter computation and updating process of the State Space Selection Mechanism in Algorithm 1. Further, the pseudo-code for the employed Selective Scan method is presented in Algorithm 2.
Algorithm 1 State Space Selection Mechanism Parameter Calculation and Upgrading.
Input: $x\in\mathbb{R}^{b\times l\times d_{\text{model}}}$, learnable model params $A_{log}\in\mathbb{R}^{d_{\text{in}}\times n}$ and $D\in\mathbb{R}^{d_{\text{in}}}$;
Output: $y\in\mathbb{R}^{b\times l\times d_{\text{inner}}}$;
1:  Compute input-independent params $A\in\mathbb{R}^{d_{\text{in}}\times n}$ and $D\in\mathbb{R}^{d_{\text{in}}}$:
2:  $A=-\exp(A_{log})$, $D=D$;
3:  Process input $x$ with Linear Layer $Input\_proj$:
4:  $x_{\text{dbl}}=Input\_proj(x)$, $x_{\text{dbl}}\in\mathbb{R}^{b\times l\times(dt_{\text{rank}}+2n)}$;
5:  Dividing $x_{\text{dbl}}$ in order to get the input-dependent params $\Delta,B,C$, where $\Delta\in\mathbb{R}^{b\times l\times dt_{\text{rank}}}$, $B,C\in\mathbb{R}^{b\times l\times n}$;
6:  Use Linear Layer $dt\_proj$ and $SoftPlus$ activation to adjust $\Delta$:
7:  $\Delta^{*}=SoftPlus(dt\_proj(\Delta))$, where $\Delta^{*}\in\mathbb{R}^{b\times l\times d_{\text{in}}}$;
8:  Employ the $SelectiveScan$ function to proceed with $x,\Delta^{*},A,B,C,D$:
9:  $y=SelectiveScan(x,\Delta^{*},A,B,C,D)$;
10:  Return the Output $y$.
Algorithm 2 Selective Scan Algorithm
Input: $u,\Delta^{*},A,B,C,D$;
Output: $z$;
1:  Obtain the dim of $u$ as $(b,l,d_{\text{in}})$ and the second dim of $A$ as $n$;
2:  Discretize continuous parameters $A$ and $B$:
3:  $\Delta A\leftarrow\exp(\text{einsum}(\Delta^{*},A))$, where $\Delta A\in\mathbb{R}^{b\times l\times d_{\text{in}}\times n}$;
4:  $\Delta B_{u}\leftarrow\text{einsum}(\Delta^{*},B,u)$, where $\Delta B_{u}\in\mathbb{R}^{b\times l\times d_{\text{in}}\times n}$;
5:  Initialize state $x$ as zeroes with dimension $\mathbb{R}^{b\times d_{\text{in}}\times n}$;
6:  for $i$ from $1$ to $l$ do
7:     $x\leftarrow\Delta A[:,i]\times x+\Delta B_{u}[:,i]$;
8:     $z\leftarrow\text{einsum}(x,C[:,i,:])$;
9:     Append $z$ to the output list $z_{s}$;
10:  end for
11:  $z\leftarrow\text{stack}(z_{s})$, where $z\in\mathbb{R}^{b\times l\times d_{\text{in}}}$;
12:  Adding the direct gain $D$:
13:  $z\leftarrow z+u\times D$
14:  Return $z$.
@@ 4.1 Dataset Statistics and Baseline Methods
Datasets. For experimental evaluation, we meticulously select three real-world STG datasets from California road network speed records (PeMS04) Song et al. (2020), Hangzhou urban metro system entry/exit records (HZMetro) Liu et al. (2022), and weather station records across mainland China (KnowAir) Wang et al. (2020). Acquired from the Caltrans Performance Measurement System (PeMS), PeMS04 records the average traffic speed from a total number of 307 sensors distributed on California Highway system. The time range of PeMS04 is 1st/Jan/2018 - 28th/Feb/2018, with 5-minute time interval. HZMetro was created with the In/Out transaction records of 80 stations of Hangzhou Metro system from 1st Jan 2019 to 25th Jan 2019, which takes 15-minute as data collection time interval. KnowAir documents weather observations in every 3-hour, encompassing a dataset that spans from 2015 to 2018. Within the KnowAir dataset, there are three subsets. In this study, we exclusively use subset No. 3, covering the period from 1/Sep/2016 to 31/Jan/2017. The detailed dataset statistics and descriptions are summarized in Table 1.
@@
STGNN-based Methods: (I) STGCN Yu et al. (2018), (II) STSGCN Song et al. (2020), (III) STG-NCDE Choi et al. (2022), (IV) DDGCRN Weng et al. (2023).
Attention (Transformer)-based Methods: (V) ASTGCN Guo et al. (2019), (VI) ASTGNN Guo et al. (2021), (VII) PDFormer Jiang et al. (2023), (VIII) STAEformer Liu et al. (2023), (IX) MultiSPANS Zou et al. (2024).
@@ 4.2 Implementation Settings
We split all the datasets with a ratio of 6:2:2 for Training/Validation/Testing sets along the time axis, respectively. Before starting model training session, all the data samples are normalized into range $[0,1]$ with MinMax normalization method. We set the batch size as 48 and employ AdamW as the Optimizer. The learning rate is initialized as $1e^{-4}$, with a weight decay of $1e^{-2}$. CosineAnnealingLR Loshchilov and Hutter (2016) is adopted as the learning rate scheduler, with a maximum of 50 iterations, and a minimum learning rate $1e^{-5}$. Model training epochs is set as 100. The number of GS3B Encoder Layer is set as $N=4$. We use the historical ground-truth data from the past 12 steps (for historical recent/period/trend scales) to forecast the future 12 time steps. Mean Squared Error (MSE) is employed as the Loss Function for optimization. And finally, we employ RMSE, MAE, MAPE as the evaluation metrics. All experiments are conducted under one NVIDIA A100 80GB GPU for neural computing acceleration.
@@ 4.5 Ablation Study
    STG-Mamba w/o KF-Upgrading: We replace the KF-Upgrading module in KFGN with the simple sum average operation to fuse the three temporal branches.
@@
    STG-Mamba w/o Dynamic Filter Linear: We replace the proposed Dynamic Filter Linear in KFGN with the basic static graph convolution.
@@
    STG-Mamba w/o SSSM: The State Space Selection Mechansim (SSSM) is removed from the ST-S3M module.
  5. 5.
    STG-Mamba w/o ST-S3M: The whole ST-S3M module is removed from GS3B Encoder.
As illustrated in Figure 5, STG-Mamba w/o KF-Upgrading shows worse performance than the full model, demonstrating the necessity and suitability of employing statistical learning-based GNN state space upgrading and optimization for SSSM-based methods. The decrease of performance due to the absence of Dynamic Filter Linear demonstrate its effectiveness, and the necessity of designing a suitable adaptive GNN for STG feature learning. Furthermore, removing the State Space Selection Mechansim in ST-S3M results in a substantial degradation of model capability, proving the feasibility of using SSSM as the substitute for Attention mechanism. Finally, the removal of ST-S3M makes the STG-Mamba degrade to a plain GNN-based model, resulting in the lowest performance.
@@ 5 Conclusions
In this work, we for the first time introduce deep learning-based selective state space models (SSSM) for spatial-temporal graph learning tasks. We propose STG-Mamba that leverages modern SSSM for accurate and efficient STG forecasting. STG-Mamba is formulated following the stacked Residual Encoder architecture, with the ST-S3M module facilitating input-dependent graph evolution and node feature selection, successfully integrates STG network with SSSMs. Through the Kalman Filtering Graph Neural Networks (KFGN), the learned STG embeddings achieve a smooth optimization & upgrading based on statistical learning, aligning with the whole STG selective state space modeling process. Compared with Attention-based methods, STG-Mamba is able to achieve linear-time complexity and substantial reduction in FLOPs and test inference time. We conduct extensive empirical studies on three open-sourced STG datasets from urban traffic, metro In/Out flow, and weather records. Experimental results demonstrate the consistent superiority of STG-Mamba over other benchmark methods. We believe that STG-Mamba presents a brand new promising approach to general STG learning fields, offering competitive model performance and input-dependent context-aware feature selection under an affordable computational cost.